# Beyond the Main Sequence: Binary Evolution Pathways to Blue Stragglers in the Gaia Era

### v1

@@ 3.5 Comparison with JS21
We recall that JS21 classified BSSs into three classes, i.e. low– (), high– () and extreme– () objects, which are speculated to predominantly form via MT, stellar collisions, and multiple-mergers/MT, respectively. Apart from our direct comparison of OCs and GCs, the difference between our results and those provided by JS21 is that we were not able to find any stars with in the extreme– regime, most likely due to our Gaia photometry cuts. It is important to mention, and as JS21 have warned, that the “fractional mass excess” does not strictly indicate that how BSSs are formed, but it rather suggests how BSSs may likely be formed. For a deeper analysis, which escapes the Gaia limitations, we do need FUV observations to be able to distinguish the presence or absence of hot companions (e.g. Subramaniam et al., 2016; Sindhu et al., 2019).
We plot the distribution of low- and high– BSSs in OCs and GCs in the Gaia CMD shown in Figure 8, similar to JS21 (see their Fig. 6). We find that BSSs in OCs and GCs classified as high– show a modest difference in color, being bluer than low– BSSs by . The more significant difference is in the absolute magnitude where high– BSSs are, on average, approximately 10 times brighter than low– BSSs. This difference is appreciable for both type of clusters, GCs and OCs. Although this empirical classification resembles the instantaneous mass excess with respect to the MSTO, it does not resemble the classic evolutionary pathways of stars, which leads them from the MS to the colder sub-giant branch and then up the RGB.
3.6 Towards BSS Evolutionary Pathways
With goal of advancing our understanding of the BSS formation channels and capturing the predominance of BSS formation channels, we combine the fractional mass excess parameter () with the our earlier derived equivalent BSS ages. We focus our attention on Figure 9, the upper panel of which reproduces the bottom right panel of Figure 5; however, this time we separate the two sequences444Do not confuse this with the BSS double sequences in a CMD discussed earlier, as the upper panel Figure 9 is not a CMD. into a high-mass excess sequence and low-mass excess through the next procedure:
@@ 3.6 Towards BSS Evolutionary Pathways
For the identified BSSs we separate them using as criteria if they belong to an OC or a GC and make a plot vs. (similar as the bottom-right panel of Figure 5).
@@
(4) where is the BSS age found with isochrone fittings. We then define the superior edge for OCs as:
(5) In a similar way, for GCs we compute these edges as:
(6) and:
@@
To check the distribution along the horizontal axis, for every point from OCs in this plane with coordinates , we compute the distance between the point and its superior edge along the vertical axis, and normalize this value by the distance between the superior and inferior edges to “flatten” it. We can reach this using the expression:
@@
We repeat the last step, but this time for BSSs from GCs using the respective edges values from step 2.
@@
If the point has a higher probability to belong to the Gaussian with lower distances, this means that the point belongs to the distribution that is closer to the superior edge and, therefore, is classified as a member of the “upper” sequence. Else, they are classified as a member of the “lower” sequence.
To properly compare BSSs with this new classification scheme, we plot them in a Gaia photometry CMD in the lower panel of Figure 9 (bottom panel), which can be directly compared to Figure 8 which encodes the classification from JS21. Our new classification now resembles more the expected stellar evolutionary pathways of BSSs, clearly separating the BSS population in GCs into two dominant sequences. We also observe that the upper- sequence BSSs in OCs populate predominantly the brighter sequence of OC BSSs in the CMD. In general, we find that the color range covered by the two sequences is similar for GCs and OCs, but there is a systematic offset in luminosity with the upper- sequence BSSs appearing about brighter than their lower- sequence counterparts.
In the following, we briefly explore possible explanations and consideration when interpreting these clearly distinct BSS sequences.
@@ 3.6.1 Mass Transfer Evolution
From a binary evolution perspective, the upper- sequence may represent a BSS population which was formed through more efficient MT processes. This would imply more accreted mass and, hence, more fuel on the receiving star, accelerating its evolution (McCrea, 1964). Assuming this scenario and a subsequent common-envelope evolution, the separation between the stars of such a binary system will eventually decrease until the final stages which could play out in two scenarios: i) the cores could merge or ii) one of the cores could be ejected from the system through interactions with envelope and/or other stars. The former scenarios is more likely (Hurley et al., 2002). The result of a stellar merger is a massive BSS, which would experience fast subsequent stellar evolution () and be more likely encountered at older BSS ages. We, therefore, interpret the two sequences in the vs. plane as post-merger/close-binary interaction sequence (“upper” sequence) and pre-merger/close-binary interaction sequence (“lower” sequence, see Fig. 9).
@@ 3.6.2 Unresolved Binary Stars
It is trivial that some BSSs may be unresolved binary systems (Gosnell et al., 2015; Subramaniam et al., 2016; Sindhu et al., 2019). However, when fitting isochrone models to our BSS sample, we are implicitly assuming single-star evolution. Jiang (2022) has shown through multiple simulations that the total mass of the system is higher in unresolved MT binaries compared with post-MT BSSs by about –. This means that the isochrone fitting method could be overestimating ages and underestimating the masses for unresolved MT BSSs, located preferentially in CMD regions with similar colors as the MSTO and the sub-giant branch. Therefore, the upper- sequence could be stars with an overestimated age –showing higher values– since they are unresolved MT binaries, while the lower- BSSs could be post-MT binaries whose companions have either negligible mass or which are in the post-MT phase.
3.6.3 The Impact of Stellar Radius Variations
As rejuvenated stars evolve of the ZAMS towards the sub-giant branch and are therefore expected to grow in size, we compute the radii of our sample BSSs, , using the relation , where is the gravitational constant, is the stellar mass and is the surface gravity. We assume in cgs units (from isochrone models) and stellar mass in units, computing the value of in solar radii using Astropy routines (Astropy Collaboration et al., 2013). We show the results in Figure 10: and as function of stellar radius (). We note that after adding stellar radius, the separation between upper- and lower- sequences is even more pronounced for GCs.
This is consistent with the expectation of an evolved binary system for BSSs in the upper- sequence (orange squares in Fig. 10), where the system might be in the common-envelope phase. Such systems evolve rapidly and reach subsequent stages with similar characteristics of an SGB or RGB star, where they begin the radial expansion, which qualitatively agrees with our observations. However, we are unable to identify any clear difference between the upper- and lower- sequences for BSSs in OCs, which may be an indication of a different mix of BSS formation mechanisms acting in lower mass, lower density star clusters. We defer the detailed modeling of these systematic differences between OCs and GCs to future work.
@@ 4 Conclusions
With the new - classification scheme of BSSs, we inspect the CMDs of all our sample GCs and OCs individually to see whether the BSS sub-populations are differentiated by our classification in the color-magnitude space. We observe that BSSs classified as upper- members are brighter whereas lower- sequence member stars appear fainter. In Figure 11 we present the distribution of upper- and lower- sequence BSSs in the representative GC NGC 362 and OC Berkeley 39. We also overlaid isochrones correspond to the ZAMS and binary-ZAMS. The figure shows that binary-ZAMS passes through majority of the BSSs intersecting both upper- and lower- sequence, while the ZAMS passes through a smaller fraction of lower- sequence (red points). The lack of BSSs on ZAMS also suggests that the lack of collisional BSSs, which is also expected since NGC 362 is a core-collapsed GC and Gaia data are unable to resolve the core of the GCs. We also do not expect to observe collisional BSSs in OCs as the environment is not dense enough to favor collision. Our classification is not similar to the literature classification of blue and red BSSs sequences observed in core-collapsed GCs, where blue BSSs are thought to form through direct collisions, and the BSSs in the red sequence form via MT (Ferraro et al., 2009; Dalessandro et al., 2013). We notice that blue (or collisional) BSSs are dominated by our lower sequence, whereas the red sequence is a mix of upper and lower sequence. Our new BSSs classification based on - mainly highlights the pre- and post- merger/MT scenarios of the BSSs. Our study suggests that the presence of companions across our - classification may imply that most of our stars classified as upper- sequence BSSs could be consistent with the presence of a companion, whereas lower- BSSs may preferentially indicate post-MT/merger products BSSs. This may support our hypothesis that assumes that we could be overestimating BSS ages due to our inability to resolve the binary system and therefore having the BSS classified as upper- sequence member.
@@
The distribution of upper- and lower- sequence BSSs in the CMDs for GCs and OCs is similar for all clusters. The CMDs of all the clusters highlighted with these two sequences are made available online for reference. In some of the clusters (NGC 104, NGC288, NGC 2808, NGC 3201, NGC 5139, NGC 6121), not classified as core-collapsed GCs, we observe a significant number of BSSs are located on the ZAMS or bluer than ZAMS, which most probably form through direct collision or post-MT/merger, as suggested by the literature on core-collapsed GCs. The presence of a significant number of BSSs outside the core or half-light radius of the clusters makes these GCs as interesting candidates and required further detailed analysis to interpret the observed nature.
@@
Much more theoretical work and observations are required to elucidate whether and to what extent our new classification scheme of BSS resembles the formation and evolution mechanisms of these unique stars. For now, we summarize in the following the results and conclusions of our work.
@@
We have selected OCs from Dias et al. (2021) that have more than 350 star members and GCs from Vasiliev & Baumgardt (2021) with more than 1000 star members to ensure well-sampled cluster stellar populations, requesting data from Gaia DR3, for which we adopt an algorithm based on prescriptions from C18 and Cordoni et al. (2020) to extract cluster members.
@@
Using the ridge line of the observed CMD and the best-fit isochrone found in the previous step along with the ZAMS and the dispersion of stars along these sequences we define a BSS selection. We are able to find BSSs in all our sample GCs and in 42 out of 129 () OCs of our sample. We identify a total of 4399 BSSs; 434 () located in OCs and 3965 () located in GCs.
@@
We extract astrophysical parameters for all identified BSSs using 3 different methods: 1) color–temperature relations (i.e. the infrared-flux method, IRFM); 2) model isochrone fitting; 3) parameters from Gaia DR3 spectra from Zhang et al. (2023, called XPP), if available. From color–temperature relations we only obtain . From isochrone fitting we obtain , , , stellar masses and ages. From XPP we obtain and .
@@
We do not find BSSs in clusters younger than .
@@
We find effective temperatures for BSSs from color–temperature relations and from isochrone–fitting for all our BSS sample. found in OC BSSs are and from color–temperature and isochrone–fitting methods, respectively. BSSs in GCs present slightly lower temperatures with derived from color–temperature relations and from isochrone models. Overall, the difference in found from isochrone–fittings are hotter than those obtained from color–temperature relations. These values are in agreement with previous studies.
@@
Based on the isochrone–fitting method, we find the average mass of all our BSSs to be, approximately, . In OCs, we find the average mass of BSSs at . For GCs we obtain an average BSS mass of . Our average mass for BSSs in OCs are lower to those found by JS21, which is likely to our more conservative selection, however, they are still within their respective error ranges. Our results are in excellent agreement with most other studies.
@@
We compare against its counterpart measured at the MSTO of the parent cluster. Independent of the method applied, for OCs and GCs we find that the difference between the mean BSS and the MSTO decreases with the cluster age. This is in line with expectations from stellar evolution because where massive stars evolve faster than stars at the MSTO. In OCs the dispersion decreases with the cluster age. We do not find this last relation in GCs, which means that the dispersion of BSSs does not depend of the cluster age in GCs.
@@
We follow the definition provided by JS21 for the “fractional mass excess”, or . This parameter quantifies the difference between the BSS mass and the stellar mass at the MSTO –this last based on isochrone-fittings–, normalized by the MSTO mass. is roughly equal/proportional to the MT efficiency in a binary system. These authors classify in three regimes, each suspected to be dominated by a specific BSS formation pathway: i) low– (, likely to be formed via MT), high– (, likely to be formed from mergers/collisions) and extreme– (, likely to be multiple-MT systems). In our sample, we do not find any extreme–, which is likely to be due to the conservative selection, where we avoided stars brighter than –to avoid Gaia systematic errors– that may contain some more massive BSSs. In our final BSS sample, is classified as high– and as low–. For OCs this proportion is and for low– and high–, respectively. For BSSs in GCs, we find these percentages to be for low– and for high–. This suggests that mass-transfer is the most likely formation channel for BSSs; but OCs show a higher fraction of high– BSSs, some of which may have formed through mergers as well.
@@
Finally, we explore a new BSS classification scheme, which uses the vs. . This parameter space shows a clear double sequence for BSS in GCs: one corresponding to an “upper” sequence, with higher as a function of the BSS age, and another “lower” sequence, linked to smaller as a function of the BSS age. We plot BSSs labeled by this classification scheme in the Gaia CMD and find a clear division between these branches; especially for GCs. BSSs on the upper- sequence are brighter than those located on the lower- sequence. We explore two possible explanations: 1) In binary star evolution, mass transfer increases the mass of the receiving star and in the process shrinks the distance between the binaries, while the receiving star begins an accelerated evolution and, hence, will appear as an “older” star in the isochrone tracks. In this scheme, these upper- and lower- sequences are consistent with post-merger/close-binary interaction and pre-merger binary evolution. 2) We might be overestimating the age and underestimating the mass of the upper- sequence BSSs, since we are assuming single star evolution during isochrone fitting. Therefore, the upper- and lower- sequences may be equivalent to an “overestimated” and “realistic” BSSs age, which could be correlated with the presence (or absence) of a BSS companion. Further observations targeting various parameters, such as the stellar rotation speed statistics, chemical compositions and binarity are rquired to shed light on this intriguing new bifurcation in the vs. classification.
All BSS parameters are available from the authors upon reasonable request.

### v2

@@ 3.5 Comparison with JS21
We recall that JS21 classified BSSs into three classes, i.e. low– (), high– () and extreme– () objects, which are speculated to predominantly form via MT, stellar collisions, and multiple-mergers/MT, respectively. In addition to our direct comparison of OCs and GCs, the difference between our results and those provided by JS21 is that we were unable to find any stars with in the extreme– regime, most likely due to our Gaia photometry cuts. It is important to mention, and as JS21 have warned, that the “fractional mass excess” does not strictly indicate how BSSs are formed, but it rather suggests how BSSs may likely be formed. For a deeper analysis, which escapes the Gaia limitations, we do need FUV observations to be able to distinguish the presence or absence of hot companions (e.g. Subramaniam et al., 2016; Sindhu et al., 2019).
We plot the distribution of low- and high– BSSs in OCs and GCs in the Gaia CMD shown in Figure 8, similar to JS21 (see their Fig. 6). We find that BSSs in OCs and GCs classified as high– show a modest difference in color, being bluer than low– BSSs by . The more significant difference is in the absolute magnitude where high– BSSs are, on average, approximately 10 times brighter than low– BSSs. This difference is appreciable for both types of clusters, GCs and OCs. Although this empirical classification resembles the instantaneous mass excess with respect to the MSTO, it does not resemble the classic evolutionary pathways of stars, which leads them from the MS to the colder subgiant branch and then up the RGB.
3.6 Toward evolutionary pathways of BSSs
With the goal of advancing our understanding of the BSS formation channels and capturing the predominance of the BSS formation channels, we combine the fractional mass excess parameter () with our earlier derived equivalent BSS ages. We focus our attention on Figure 9, the upper panel of which reproduces the bottom right panel of Figure 5, but this time, we separate the two sequences555Do not confuse this with the BSS double sequences in a CMD discussed earlier, as the upper panel Figure 9 is not a CMD. into a high-mass excess sequence and low-mass excess through the next procedure:
@@ 3.6 Toward evolutionary pathways of BSSs
For the identified BSSs, we separate them using as criteria if they belong to an OC or a GC and make a plot versus (similar as the bottom-right panel of Figure 5).
@@
(4) where is the BSS age found with isochrone fittings. We then define the superior edge for OCs as
(5) In a similar way, we computed these edges for GCs as
(6) and
@@
To check the distribution along the horizontal axis, for every point from OCs in this plane with coordinates , we compute the distance between the point and its superior edge along the vertical axis and normalize this value by the distance between the superior and inferior edges to “flatten” it. We can reach this using the expression
@@
We repeated the last step, but this time for BSSs from GCs using the respective edges values from step 2.
@@
If the point has a higher probability to belong to the Gaussian with lower distances, this means that the point belongs to the distribution that is closer to the superior edge, and it is therefore classified as a member of the “upper” sequence. Else, they are classified as a member of the “lower” sequence.
To properly compare BSSs with this new classification scheme, we plot them in a Gaia photometry CMD in the lower panel of Figure 9 (bottom panel), which can be directly compared to Figure 8 which encodes the classification from JS21. Our new classification now resembles more the expected stellar evolutionary pathways of BSSs, clearly separating the BSS population in GCs into two dominant sequences. We also observe that the upper- sequence BSSs in OCs predominantly populate the brighter sequence of OC BSSs in the CMD. In general, we find that the color range covered by the two sequences is similar for GCs and OCs, but there is a systematic offset in luminosity with the upper- sequence BSSs appearing about brighter than their lower- sequence counterparts.
In the following, we briefly explore possible explanations and considerations when interpreting these clearly distinct BSS sequences.
@@ 3.6.1 Mass transfer evolution
From a binary evolution perspective, the upper- sequence may represent a BSS population that was formed through more efficient MT processes. This would imply more accreted mass and, hence, more fuel on the receiving star, accelerating its evolution (McCrea, 1964). Assuming this scenario and a subsequent common-envelope evolution, the separation between the stars of such a binary system will eventually decrease until the final stages which could play out in two scenarios: i) the cores could merge or ii) one of the cores could be ejected from the system through interactions with envelope and/or other stars. The former scenario is more likely (Hurley et al., 2002). The result of a stellar merger is a massive BSS, which would experience fast subsequent stellar evolution () and be more likely encountered at older BSS ages. We therefore interpret the two sequences in the versus plane as post-merger/close-binary interaction sequence (“upper” sequence) and pre-merger/close-binary interaction sequence (“lower” sequence, see Fig. 9).
@@ 3.6.2 Unresolved binary stars
It is trivial that some BSSs may be unresolved binary systems (Gosnell et al., 2015; Subramaniam et al., 2016; Sindhu et al., 2019). When we fit isochrone models to our BSS sample, however, we implicitly assumed a single-star evolution. Jiang (2022) has shown through multiple simulations that the total mass of the system is higher in unresolved MT binaries compared to post-MT BSSs by about –. This means that the isochrone fitting method could be overestimating ages and underestimating the masses for unresolved MT BSSs, located preferentially in CMD regions with similar colors as the MSTO and the subgiant branch. Therefore, the upper- sequence could be stars with an overestimated age –showing higher values– since they are unresolved MT binaries, while the lower- BSSs could be post-MT binaries whose companions either have negligible mass or which are in the post-MT phase.
3.6.3 Effect of variations in the stellar radius
As rejuvenated stars evolve of the ZAMS towards the subgiant branch and are therefore expected to grow in size, we compute the radii of our sample BSSs, , using the relation , where is the gravitational constant, is the stellar mass and is the surface gravity. We assume in cgs units (from isochrone models) and stellar mass in units, computing the value of in solar radii using Astropy routines (Astropy Collaboration et al., 2013). We show the results in Figure 10: and as a function of stellar radius (). We note that after adding stellar radius, the separation between upper- and lower- sequences is even more pronounced for GCs.
This is consistent with the expectation of an evolved binary system for BSSs in the upper- sequence (orange squares in Fig. 10), where the system might be in the common envelope phase. Such systems evolve rapidly and reach subsequent stages with similar characteristics of an SGB or RGB star, where they begin the radial expansion, which qualitatively agrees with our observations. We were unable to identify any clear difference between the upper- and lower- sequences for BSSs in OCs, however, which may be an indication of a different mix of BSS formation mechanisms acting in lower mass, lower density star clusters. We defer the detailed modeling of these systematic differences between OCs and GCs to future work.
@@ 4 Conclusions
With the new - classification scheme of BSSs, we inspected the CMDs of all our sample GCs and OCs individually to determine whether the BSS subpopulations are differentiated by our classification in the color-magnitude space. We observed that BSSs classified as upper- members are brighter and that lower- sequence member stars appear fainter. In Figure 11 we presented the distribution of upper- and lower- sequence BSSs in the representative GC NGC 362 and OC Berkeley 39. We also overlaid isochrones corresponding to the ZAMS and binary ZAMS. The figure shows that binary ZAMS passes through the majority of the BSSs that intersect the upper- and lower- sequence, while the ZAMS passes through a smaller fraction of the lower- sequence (red points). The lack of BSSs on the ZAMS also suggests the lack of collisional BSSs, which is also expected because NGC 362 is a core-collapse GC and the Gaia data are unable to resolve the core of the GCs. We do not expect to observe collisional BSSs in OCs either because the environment is not dense enough to favor collision. Our classification is not similar to the literature classification of blue and red BSSs sequences observed in core-collapse GCs, where blue BSSs are thought to form through direct collisions and the BSSs in the red sequence are thought to form via MT (Ferraro et al., 2009; Dalessandro et al., 2013). We note that blue (or collisional) BSSs are dominated by our lower sequence, whereas the red sequence is a mix of the upper and lower sequence. Our new BSSs classification based on - mainly highlights the pre- and post- merger/MT scenarios of the BSSs. Our study suggests that companions in our - classification may imply that most of our stars that are classified as upper- sequence BSSs might be consistent with the presence of a companion, while lower- BSSs may preferentially indicate post-MT/merger-product BSSs. This may support our hypothesis, which assumes that we might be overestimating the ages of BSSs because we cannot resolve the binary system, and therefore, the BSSs were classified as upper- sequence members.
@@
The distribution of upper- and lower- sequence BSSs in the CMDs for GCs and OCs is similar for all clusters. The CMDs of all the clusters highlighted with these two sequences are made available on Zenodo for reference. In some of the clusters (NGC 104, NGC288, NGC 2808, NGC 3201, NGC 5139, and NGC 6121) that are not classified as core-collapse GCs, we observed that a significant number of BSSs are located on the ZAMS or are bluer than the ZAMS. These most probably formed through a direct collision or post-MT/merger, as suggested by the literature on core-collapse GCs. The significant number of BSSs outside the core or half-light radius of the clusters makes these GCs interesting candidates and requires a further detailed analysis to interpret the observed nature.
@@
Much more theoretical work and observations are required to elucidate whether and to which extent our new classification scheme of BSSs resembles the formation and evolution mechanisms of these unique stars. We summarize the results and conclusions of our work below.
@@
We selected OCs from Dias et al. (2021) with more than 350 star members and GCs from Vasiliev & Baumgardt (2021) with more than 1000 star members to ensure well-sampled cluster stellar populations, and we used data from Gaia DR3, for which we adopted an algorithm based on prescriptions from C18 and Cordoni et al. (2020) to extract cluster members.
@@
Using the ridge line of the observed CMD and the best-fit isochrone found in the previous step along with the ZAMS and the dispersion of stars along these sequences, we defined a BSS selection. We were able to find BSSs in all our sample GCs and in 42 out of 129 () OCs of our sample. We identified a total of 4399 BSSs, 434 of which () are located in OCs and 3965 () are located in GCs.
@@
We extracted astrophysical parameters for all identified BSSs using three different methods: 1) color–temperature relations (i.e., IRFM), 2) model isochrone fitting, and 3) parameters from Gaia DR3 spectra from Zhang et al. (2023, called XPP), if available. From the color–temperature relations, we only obtained . From the isochrone fitting, we obtained , , , stellar masses, and ages. From XPP we obtained and .
@@
We found no BSSs in clusters younger than .
@@
We found effective temperatures for BSSs from the color–temperature relations and from the isochrone fitting for our entire BSS sample. The found in OC BSSs are and from the color–temperature and isochrone–fitting methods, respectively. The BSSs in GCs present slightly lower temperatures, with derived from color–temperature relations and from isochrone models. Overall, the difference in found from isochrone fittings is that they are hotter than those obtained from color–temperature relations. These values agree with those of previous studies.
@@
Based on the isochrone–fitting method, we determined that the average mass of all our BSSs is approximately . The average mass of BSSs in OCs is at . We obtained an average BSS mass for GCs of . Our average mass for BSSs in OCs is lower than the mass determined by JS21, which is likely due to our more conservative selection, but they still agree within their respective error ranges. Our results agree very well with most other studies.
@@
We compared against its counterpart measured at the MSTO of the parent cluster. Regardless of the method we applied, we found for OCs and GCs that the difference between the mean BSS and the MSTO decreases with the cluster age. This is in line with expectations from stellar evolution because massive stars evolve faster than stars at the MSTO. The dispersion in OCs decreases with the cluster age. We did not find this last relation in GCs, which means that the dispersion of BSSs does not depend on the cluster age in GCs.
@@
We followed the definition provided by JS21 for the fractional mass excess, or . This parameter quantifies the difference between the BSS mass and the stellar mass at the MSTO (this last based on isochrone fittings), normalized by the MSTO mass. is roughly equal or proportional to the MT efficiency in a binary system. These authors classified into three regimes, each assumed to be dominated by a specific BSS formation pathway: i) low– (, likely to be formed via MT), high– (, likely to be formed through mergers or collisions), and extreme– (, likely to be multiple MT systems). In our sample, we found no extreme–, which is likely to be due to the conservative selection, where we avoided stars brighter than (to avoid Gaia systematic errors), which may contain some more massive BSSs. In our final BSS sample, are classified as high– and as low–. This proportion is and for low– and high–, respectively, for OCs. For BSSs in GCs, we found these percentages to be for low– and for high–. This suggests that mass transfer is the most likely formation channel for BSSs, but OCs show a higher fraction of high– BSSs, some of which may have formed through mergers as well.
@@
Finally, we explored a new BSS classification scheme that uses versus . This parameter space shows a clear double sequence for BSS in GCs: One parameter space corresponds to an upper sequence, with higher as a function of the BSS age, and another lower sequence that is linked to lower as a function of the BSS age. We plotted the BSSs labeled according to this classification scheme in the Gaia CMD and found a clear division between these branches, especially for GCs. BSSs on the upper- sequence are brighter than those located on the lower- sequence. We explored two possible explanations: 1) In binary star evolution, mass transfer increases the mass of the receiving star and in the process decreases the distance between the binaries, while the receiving star begins an accelerated evolution and therefore appear as an older star in the isochrone tracks. In this scheme, these upper- and lower- sequences are consistent with post-merger or close-binary interaction and pre-merger binary evolution. 2) We might be overestimating the age and underestimating the mass of the upper- sequence BSSs because we assumed single-star evolution during the isochrone fitting. Therefore, the upper- and lower- sequences may be equivalent to an overestimated and realistic BSSs age, which could be correlated with the presence (or absence) of a BSS companion. Further observations targeting various parameters, such as the stellar rotation speed statistics, chemical compositions, and binarity, are required to shed light on this intriguing new bifurcation in the versus classification.
5 Data availability
Parameters of BSSs, YSSs, RSSs and the star clusters presented in this study (full Tables D.1 and D.2) are available at the CDS via https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/699/A142 and on Zenodo for reference. The CMDs for all the clusters as shown in Figure 11 are also made available on Zenodo.